# Detecting Shearless Phase-Space Transport Barriers in Global Gyrokinetic Turbulence Simulations with Test Particle Map Models

### v1

@@ Abstract
In magnetically confined fusion plasmas, the role played by zonal $E\times B$ flow shear layers in the suppression of turbulent transport is relatively well-understood. However, less is understood about the role played by the weak shear regions that arise in the non-monotonic radial electric field profiles often associated with these shear layers. In electrostatic simulations from the global total-$f$ gyrokinetic particle-in-cell code XGC, we demonstrate how shearless regions with non-zero flow curvature form zonal “jets” that, in conjunction with neighboring regions of shear, can act as robust barriers to particle transport and turbulence spreading. By isolating quasi-coherent fluctuations radially localized to the zonal jets, we construct a map model for the Lagrangian dynamics of gyrokinetic test particles in the presence of drift waves. We identify the presence of shearless invariant tori in this model and verify that these tori act as partial phase-space transport barriers in the simulations. We also demonstrate how avalanches impinging on these shearless tori cause reconnection events that form “cold/warm core ring” structures analogous to those found in oceanic jets, facilitating transport across the barriers without destroying them completely. We discuss how shearless tori may generically arise from tertiary instabilities or other types of discrete eigenmodes, suggesting their potential relevance to broader classes of turbulent fluctuations.
@@ 1 Introduction
Sheared $E\times B$ flows arising from the toroidally symmetric component of the radial electric field $E_{r}$, also known as zonal flows, are central to turbulence regulation in tokamak plasmas (Diamond et al., 2005; Burrell, 2020; Staebler et al., 2024). A key thrust in magnetic confinement fusion is thus to understand the mechanisms which determine the coupled intensities of zonal flows and turbulent fluctuations, which in turn predict the overall level of cross-field turbulent transport. In the core, zonal flows are known to play a key role the nonlinear upshift of the critical gradient for the onset of turbulent transport, known as the “Dimits shift” (Dimits et al., 2000). In the H-mode edge and other edge transport barrier regimes, the sheared flows are strong enough to completely quench or otherwise severely limit the amount of transport due to turbulence, leading to enhanced plasma confinement. In the L-mode edge, the growth of sheared zonal flows likely plays a key role in triggering the L-H transition (Kim & Diamond, 2003; Schmitz et al., 2012), while the collapse of sheared flows in the plasma edge has been proposed as a mechanism for setting the L-mode density limit (Singh & Diamond, 2021; Diamond et al., 2023).
@@
There is a well-developed understanding of how local $E\times B$ shear can lead to the saturation or suppression of turbulence (Biglari et al., 1990; Hahm & Burrell, 1995; Waltz et al., 1998). Work in eikonal theory and wave kinetics has developed a detailed picture of how the radial variation of the zonal shear can affect drift wave turbulence in the scale-separated geometric optics level (Smolyakov et al., 2000; Ruiz et al., 2016). However, recent work has pointed out the role that effects beyond the geometric optics level play in drift waves-zonal flow interaction, in particular the theory of the tertiary instability (Zhu et al., 2020a, b). Quantization of the tertiary instability modes can lead to a radial localization of the drift-wave envelopes which can stabilize instabilities at gradients above the critical gradient of the primary linear instability, which has been indicated as a potential mechanism for the nonlinear critical gradient associated with the Dimits shift (Kobayashi & Rogers, 2012). Shearless regions play a key role in tertiary instablity theory, as they are associated with the X- and O-points in the drifton phase space (Sasaki et al., 2017; Zhu et al., 2018) where tertiary instabilities can localize around.
@@
The central theme of this work is to show that the localization of fluctuations to shearless regions can counterintuitively assist in suppression of turbulent transport through the formation of shearless transport barriers. In the dynamical systems literature, shearless transport barriers refer to a particular class of invariant tori present in Hamiltonian flows and maps (del Castillo-Negrete, 2000; Morrison, 2000; Caldas et al., 2012), and in this context are also known as shearless or nontwist tori. Here, we use the terminology ‘shearless tori’ to refer to invariant tori in collisionless test particle dynamics, ‘shearless transport barriers’ to refer to their macroscopic manifestation in turbulent transport, and ‘shearless phase-space transport barriers’ to refer to their manifestation in particle phase space. In fixed-parameter planar maps or periodic planar flows invariant tori act as complete barriers to transport, but such conclusions do not immediately apply when considering transport due to non-autonomous perturbations, higher-dimensional phase space, and collisional effects experienced by particles in gyrokinetic turbulence.
@@
The theory of shearless transport barriers originally arose to explain observations of robust barriers to the transport of dye across zonal jets in rotating tank experiments (Behringer et al., 1991; del Castillo-Negrete & Morrison, 1993). Shearless tori have been extensively studied in a number of different systems, where they have been observed phenomenologically to be robust barriers to transport. These systems include 2-dimensional models (del Castillo-Negrete et al., 1996, 1997; Balescu, 1998; del Castillo-Negrete, 2000; Marcus et al., 2008; Da Fonseca et al., 2014), 3-dimensional models including magnetic shear and $E\times B$ shear (Horton et al., 1998; Marcus et al., 2019; Osorio-Quiroga et al., 2023; Grime et al., 2023), and in gyrokinetic models with prescribed perturbations (Anastassiou et al., 2024). There also exist generalizations of the notion of shearless transport barriers to finite-time systems (Rypina et al., 2007a; Beron-Vera et al., 2010; Farazmand et al., 2014; Falessi et al., 2015).
@@
We also consider the dynamics of the shearless phase-space transport barriers in fully developed turbulence beyond the Poincaré map model. Using tools from topological data analysis (TDA), we demonstrate a correlation between heat flux avalanche events observed in the turbulence along with the radial propagation of phase space hole/blob features in the XGC simulations. Upon reaching the shearless region, we show that these blobs cause reconnection events on the shearless phase-space transport barrier, facilitating transport across the barrier without entirely destroying it. We point out an analogy between this process and the formation of “warm core ring” and “cold core ring” structures that result from eddy detachment in oceanic jets (The Ring Group, 1981; Olson, 1991).
@@ 2.1 XGC Simulations
This work considers turbulence simulations carried out using the global total-$f$ gyrokinetic particle-in-cell code XGC1 (Ku et al., 2018b, a; Hager et al., 2022). The results have been originally reported in Zhu et al. (2024a), where the simulation data is also provided (Zhu et al., 2024b). The simulations primarily model the interactions between ion temperature gradient (ITG) turbulence and zonal flows during an early ELM-free H-mode using realistic DIII-D tokamak geometry. The equilibrium profiles are adapted from DIII-D shot number 141451 (Müller et al., 2011a, b).
@@
On the right-hand side, $C_{s}$ is a multi-species Fokker-Planck-Landau collision operator (Yoon & Chang, 2014; Hager et al., 2016), $S_{s}$ describes heating, and $N_{s}$ describes neutral ionization and charge exchange (Ku et al., 2018a). In this simulation, turbulence is dominated by ion dynamics driven by the ion temperature gradient, and a 1MW heating is applied to ions in the core to sustain the temperature gradient. Neutral dynamics are also included in the edge and scrape-off layer, providing the only particle source in the simulations, as there was no significant beam fueling. The simulation consists of 16 poloidal planes that span $1/3$ of the torus, with approximately 132k mesh nodes per poloidal plane. We refer the reader to Zhu et al. (2024a) for more details of the simulation.
Figure 1: (a-b) Equilibrium profiles used to initialize the simulations. The region of interest is marked between the vertical gray lines. (c) Poloidal cross-section showing the non-zonal component of the electrostatic potential, as as $v_{E}$ on the outboard midplane. (d-h) Temporally-averaged quantities after several turbulence times as a function of radial coordinate, including the zonal $E\times B$ rotation rate $\Omega_{E}$, the Waltz-Miller shearing rate $\gamma_{E}$, ion gyrocenter density gradient scale length $a/L_{N_{i}}$, ion temperature gradient $a/L_{T_{i}}$, and density skewness $\langle\tilde{n}_{e}^{3}\rangle/\langle\tilde{n}_{e}^{2}\rangle^{3/2}$. The shearless region $\psi=\psi^{*}$ is marked with a dashed vertical line.
@@ 2.2 Properties of the Shearless Region
From panel (g), the shearless region also appears to separate the region of flat $a/L_{T_{i}}$ in the inner core from the region of steeper $a/L_{T_{i}}$ in the edge. This is reminiscent of the argument in Singh & Diamond (2020) which argues that turbulence spreading from the edge contributes to the weakening of temperature stiffness in NML. To quantify the interaction between the shearless region and turbulence propagation, we consider the (normalized) density skewness $\langle\tilde{n}_{e}^{3}\rangle/\langle\tilde{n}_{e}^{2}\rangle^{3/2}$. Skewness measures the asymmetry of a probability distribution about its mean, with positive values indicating an excess of positive fluctuations and a negative value indicating an excess of negative fluctuations. When these fluctuations are carried by isolated structures, they are often referred to as density “blobs” or “holes” respectively. Density skewness is frequently used as an indicator of blob/hole and avalanche dynamics in both experimental and theoretical studies (D’Ippolito et al., 2011). We will show later in §4.2 that these statistical fluctuations are associated with filamentary phase-space structures, so we will also refer to them as blobs/holes. Panel (h) shows that the zero of the density skewness, associated with a transition from hole- to blob-dominated fluctuations, occurs at the same radial location as the shearless region. This suggests that the shearless region acts as the boundary between the inner core and the edge, preventing turbulence spreading from one region to the other.
@@
Moving to the time-dependent behavior of the zonal jet, in Figure 2, we show a sequence of color plots showing the evolution of various flux-surface-averaged quantities over the time and radial coordinates. Panels (a,b) illustrate the zonal jet and the zonal shear parameter $\gamma_{E}$. There is a persistent region of zero shear, equivalently a local minimum in $\Omega_{E}$, in the vicinity of the zonal jet near $\psi_{n}\approx 0.8$. Outside of the zonal jet, the zonal flows display more fluctuating behavior, with a ‘fishbone’ behavior in the space-time plots that corresponds to radial propagation. Notice that these fishbone fluctuations converge towards and terminate at the zonal jet.
These fishbone fluctuations can be correlated with the radial propagation of turbulence avalanches. In panel (c) we show the normalized flux-surface-averaged density fluctuations $\langle\tilde{n}_{e}^{2}\rangle^{1/2}/\langle n_{0}\rangle$. The density fluctuations also display fishbone-like behavior, with bursts of fluctuation amplitude forming in the inner core and the edge before converging towards the zonal jet. This behavior is also reflected in the radial energy flux $\langle\mathcal{E}\dot{\mathbf{R}}\cdot\nabla\psi\rangle$, shown in panel (d), where $\mathcal{E}$ is the particle kinetic energy and $\dot{\mathbf{R}}$ is the particle drift. Panel (e) shows the density skewness, which again shows fishbone-like behavior indicative of blobs/holes forming then converging towards the zonal jet. Similar behavior is observed in the ion temperature skewness $\langle\tilde{T}_{i}^{3}\rangle/\langle\tilde{T}_{i}^{2}\rangle^{3/2}$, which is shown in panel (f).
@@ 3.1 Construction of the Model Dynamical System
Such modes correspond to rigidly toroidally rotating fluctuations, with a toroidal angular phase velocity of $\Omega=\omega/n$. Hamiltonians consisting of rigidly toroidally rotating modes with a common angular phase velocity $\Omega$ will satisfy $\partial_{t}H+\Omega\partial_{\varphi}H=0$. Particles undergoing motion in such fields will have two invariants of motion, $(\mu,K)$, where $K=H-\Omega P_{\varphi}$ is the Hamiltonian in the rotating frame. This fact is often utilized in studies of energetic particle transport (Hsu & Sigmar, 1992; Todo, 2019). Note that a regime where most fluctuations share a nearly common angular phase velocity of $\Omega$ would correspond to weakly dispersive turbulence.
The key hypothesis we take is that in the presence of a zonal jet, the phenomenon of wave trapping will produce fluctuations which lead to this weakly dispersive regime. In Zhu et al. (2020a, b), it was demonstrated how wave trapping and anti-trapping play a key role determining the Dimits shift regime in fluid models of drift-wave turbulence. The key physics of wave trapping is that waves can constructively interfere. Physically, one can imagine that the $E\times B$ jet acts as a “cavity” which modifies the structure of the primary ITG instability. In the presence of shear of opposing signs, discrete quantum harmonic oscillator-like eigenmodes can appear which have radially-localized fluctuation envelopes. Intuitively, this radial localization increases the effective radial wavenumber, typically leading to a stabilizing effect on radial-gradient-driven instabilities. In such a regime turbulent mixing can become strongly spatially inhomogeneous; see for example Cao & Qi (2023, 2024) for detailed studies of the nonlinear dynamics of large-amplitude tertiary instabilities in a variant of the Hasegawa-Wakatani equations.
Figure 3: (a) Toroidal mode spectra and (b) angular phase velocities, both measured at one instant in time. $\langle\phi_{n}^{2}\rangle$ is the flux surface averaged squared electrostatic potential, and $\hat{\Omega}_{n}$ is the toroidally-directed angular phase velocity. The quantity $\Gamma_{0}k_{\theta}^{2}\langle\phi_{n}^{2}\rangle$ mimics the gyroaveraged $\langle\mathcal{J}[E_{\perp}]^{2}\rangle$ spectrum. Different toroidal mode numbers are shown with different colors and linestyles. $\Omega_{E}$ is shown on both plots for reference, and a dashed black line is also plotted over the angular phase velocities show the rotation rate $\Omega$ used for the model fluctuations.
@@
Figure 4: Poloidal slices of (a) The electrostatic potential $\phi_{n}$ for the $n=39$ Fourier mode, and (b) the single-mode model electrostatic potential $\delta\hat{\psi}$ at a single instant in time. The fields are plotted against the radial coordinate $\psi_{n}$ and the perpendicular field line label $\alpha=\varphi-q\theta$. The shearless region $\psi=\psi^{*}$ is demarcated with a dashed gray line. (c) Space-time plot showing the value of the electrostatic potential fluctuations $\delta\phi$, evaluated at the shearless region $\psi=\psi^{*}$ plotted against $\alpha$. A line with $\varphi\propto\Omega t$ is shown demonstrating the fixed phase velocity of the fluctuations.

### v2

@@ Abstract
In magnetically confined fusion plasmas, the role played by zonal $E\times B$ flow shear layers in the suppression of turbulent transport is relatively well-understood. However, less is understood about the role played by the weak shear regions that arise in the non-monotonic radial electric field profiles often associated with these shear layers. In electrostatic simulations from the global total-$f$ gyrokinetic particle-in-cell code XGC, we demonstrate how shearless regions with non-zero flow curvature form zonal “jets” that, in conjunction with neighboring regions of shear, can act as robust barriers to particle transport and turbulence spreading. By isolating quasi-coherent fluctuations radially localized to the zonal jets, we construct a map model for the Lagrangian dynamics of gyrokinetic test particles in the presence of drift waves. We identify the presence of shearless invariant tori in this model and verify that these tori act as partial phase-space transport barriers in the simulations. We also demonstrate how avalanches impinging on these shearless tori cause eddy detachment events that form “cold/warm core ring” structures analogous to those found in oceanic jets, facilitating transport across the barriers without destroying them completely. We discuss how shearless tori may generically arise from tertiary instabilities or other types of discrete eigenmodes, suggesting their potential relevance to broader classes of turbulent fluctuations.
@@ 1 Introduction
Sheared $E\times B$ flows arising from the toroidally symmetric component of the radial electric field $E_{r}$, also known as zonal flows, are central to turbulence regulation in tokamak plasmas (Diamond et al., 2005; Burrell, 2020; Staebler et al., 2024). A key thrust in magnetic confinement fusion is thus to understand the mechanisms which determine the coupled intensities of zonal flows and turbulent fluctuations, which in turn predict the overall level of cross-field turbulent transport. In the core, zonal flows are known to play a key role the nonlinear upshift of the critical gradient for the onset of turbulent transport, known as the “Dimits shift” (Dimits et al., 2000). In the H-mode edge and other edge transport barrier regimes, the sheared flows are strong enough to completely quench or otherwise severely limit the amount of transport due to turbulence, leading to enhanced plasma confinement. In the L-mode edge, the growth of sheared zonal flows likely plays a key role in triggering the L-H transition (Kim and Diamond, 2003; Schmitz et al., 2012), while the collapse of sheared flows in the plasma edge has been proposed as a mechanism for setting the L-mode density limit (Singh and Diamond, 2021; Diamond et al., 2023).
@@
There is a well-developed understanding of how local $E\times B$ shear can lead to the saturation or suppression of turbulence (Biglari et al., 1990; Hahm and Burrell, 1995; Waltz et al., 1998). Work in eikonal theory and wave kinetics has developed a detailed picture of how the radial variation of the zonal shear can affect drift wave turbulence in the scale-separated geometric optics level (Smolyakov et al., 2000; Ruiz et al., 2016). However, recent work has pointed out the role that effects beyond the geometric optics level play in drift waves-zonal flow interaction, in particular the theory of the tertiary instability (Zhu et al., 2020a, b). Quantization of the tertiary instability modes can lead to a radial localization of the drift-wave envelopes which can stabilize instabilities at gradients above the critical gradient of the primary linear instability, which has been indicated as a potential mechanism for the nonlinear critical gradient associated with the Dimits shift (Kobayashi and Rogers, 2012). Shearless regions play a key role in tertiary instablity theory, as they are associated with the X- and O-points in the drifton phase space (Sasaki et al., 2017; Zhu et al., 2018) where tertiary instabilities can localize around.
@@
The central theme of this work is to show that the localization of fluctuations to shearless regions can counterintuitively assist in suppression of turbulent transport through the formation of shearless transport barriers. In the dynamical systems literature, shearless transport barriers refer to a particular class of invariant tori present in Hamiltonian flows and maps (del-Castillo-Negrete, 2000; Morrison, 2000; Caldas et al., 2012), and in this context are also known as shearless or nontwist tori. Here, we use the terminology ‘shearless tori’ to refer to invariant tori in collisionless test particle dynamics, ‘shearless transport barriers’ to refer to their macroscopic manifestation in turbulent transport, and ‘shearless phase-space transport barriers’ to refer to their manifestation in particle phase space. In fixed-parameter planar maps or periodic planar flows invariant tori act as complete barriers to transport, but such conclusions do not immediately apply when considering transport due to non-autonomous perturbations, higher-dimensional phase space, and collisional effects experienced by particles in gyrokinetic turbulence.
@@
The theory of shearless transport barriers originally arose to explain observations of robust barriers to the transport of dye across zonal jets in rotating tank experiments (Behringer et al., 1991; del-Castillo-Negrete and Morrison, 1993). Shearless tori have been extensively studied in a number of different systems, where they have been observed phenomenologically to be robust barriers to transport. These systems include 2-dimensional models (del-Castillo-Negrete et al., 1996, 1997; Balescu, 1998; del-Castillo-Negrete, 2000; Marcus et al., 2008; Da Fonseca et al., 2014), 3-dimensional models including magnetic shear and $E\times B$ shear (Horton et al., 1998; Marcus et al., 2019; Osorio-Quiroga et al., 2023; Grime et al., 2023), and in gyrokinetic models with prescribed perturbations (Anastassiou et al., 2024). There also exist generalizations of the notion of shearless transport barriers to finite-time systems (Rypina et al., 2007a; Beron-Vera et al., 2010; Farazmand et al., 2014; Falessi et al., 2015).
@@
We also consider the dynamics of the shearless phase-space transport barriers in fully developed turbulence beyond the Poincaré map model. Using tools from topological data analysis (TDA), we demonstrate a correlation between heat flux avalanche events observed in the turbulence along with the radial propagation of phase space hole/blob features in the XGC simulations. Upon reaching the shearless region, we show that these blobs cause eddy detachment events associated with the zonal jet, facilitating transport across the barrier without entirely destroying it. We point out an analogy between this process and the formation of “warm core ring” and “cold core ring” structures in oceanic jets (The Ring Group, 1981; Olson, 1991).
@@ 2.1 XGC Simulations
This work considers turbulence simulations carried out using the global total-$f$ gyrokinetic particle-in-cell code XGC1 (Ku et al., 2018b, a; Hager et al., 2022). The results have been originally reported in Zhu et al. (2024b), where the simulation data is also provided (Zhu et al., 2024a). The simulations primarily model the interactions between ion temperature gradient (ITG) turbulence and zonal flows during an early ELM-free H-mode using realistic DIII-D tokamak geometry. The equilibrium profiles are adapted from DIII-D shot number 141451 (Müller et al., 2011b, a).
@@
On the right-hand side, $C_{s}$ is a multi-species Fokker-Planck-Landau collision operator (Yoon and Chang, 2014; Hager et al., 2016), $S_{s}$ describes heating, and $N_{s}$ describes neutral ionization and charge exchange (Ku et al., 2018a). In this simulation, turbulence is dominated by ion dynamics driven by the ion temperature gradient, and a 1MW heating is applied to ions in the core to sustain the temperature gradient. Neutral dynamics are also included in the edge and scrape-off layer, providing the only particle source in the simulations, as there was no significant beam fueling. The simulation consists of 16 poloidal planes that span $1/3$ of the torus, with approximately 132k mesh nodes per poloidal plane. We refer the reader to Zhu et al. (2024b) for more details of the simulation.
Figure 1: (a-b) Equilibrium profiles used to initialize the simulations. The region of interest is marked between the vertical gray lines. (c) A poloidal cross-section showing the non-zonal component of the electrostatic potential $\delta\phi$, as well as the $E\times B$ velocity $v_{E}$ evaluated at the outboard midplane. The region of interest is between the flux surfaces indicated by gray contours. (d-h) Temporally-averaged quantities after several turbulence times as a function of radial coordinate, including the zonal $E\times B$ rotation rate $\Omega_{E}$, the Waltz-Miller shearing rate $\gamma_{E}$, ion gyrocenter density gradient scale length $a/L_{N_{i}}$, ion temperature gradient $a/L_{T_{i}}$, and density skewness $\langle\tilde{n}_{e}^{3}\rangle/\langle\tilde{n}_{e}^{2}\rangle^{3/2}$. The shearless region $\psi=\psi^{*}$ is marked with a dashed vertical line.
@@ 2.2 Properties of the Shearless Region
From panel (g), the shearless region also appears to separate the region of flat $a/L_{T_{i}}$ in the inner core from the region of steeper $a/L_{T_{i}}$ in the edge. This is reminiscent of the argument in Singh and Diamond (2020) which argues that turbulence spreading from the edge contributes to the weakening of temperature stiffness in NML. To quantify the interaction between the shearless region and turbulence propagation, we consider the (normalized) density skewness $\langle\tilde{n}_{e}^{3}\rangle/\langle\tilde{n}_{e}^{2}\rangle^{3/2}$. Skewness measures the asymmetry of a probability distribution about its mean, with positive values indicating an excess of positive fluctuations and a negative value indicating an excess of negative fluctuations. When these fluctuations are carried by isolated structures, they are often referred to as density “blobs” or “holes” respectively. Density skewness is frequently used as an indicator of blob/hole and avalanche dynamics in both experimental and theoretical studies (D’Ippolito et al., 2011). We will show later in §4.2 that these statistical fluctuations are associated with filamentary phase-space structures, so we will also refer to them as blobs/holes. Panel (h) shows that the zero of the density skewness, associated with a transition from hole- to blob-dominated fluctuations, occurs at the same radial location as the shearless region. This suggests that the shearless region acts as the boundary between the inner core and the edge, preventing turbulence spreading from one region to the other.
@@
Moving to the time-dependent behavior of the zonal jet, in Figure 2, we show a sequence of color plots showing the evolution of various flux-surface-averaged quantities over the time and radial coordinates. Panels (a,b) illustrate the zonal jet and the zonal shear parameter $\gamma_{E}$. There is a persistent region of zero shear, equivalently a local minimum in $\Omega_{E}$, in the vicinity of the zonal jet near $\psi_{n}\approx 0.8$. Outside of the zonal jet, the zonal flows display more fluctuating behavior. The space-time plots suggest intermittent increases in zonal flow strength that begin in the inner core and edge which then converge radially to the zonal jet.
These radially converging zonal flow bursts can be correlated with the radial propagation of turbulence avalanches. In panel (c) we show the normalized flux-surface-averaged density fluctuations $\langle\tilde{n}_{e}^{2}\rangle^{1/2}/\langle n_{0}\rangle$. The density fluctuations also display this radial convergence behavior, with bursts of fluctuation amplitude that form in the inner core and the edge and converge towards the zonal jet. This behavior is also reflected in the radial energy flux $\langle\mathcal{E}\dot{\mathbf{R}}\cdot\nabla\psi\rangle$, shown in panel (d), where $\mathcal{E}$ is the particle kinetic energy and $\dot{\mathbf{R}}$ is the particle drift. Panel (e) shows the density skewness, which again shows this radial converging behavior indicative of blobs/holes forming then propagating towards the zonal jet. Similar behavior is observed in the ion temperature skewness $\langle\tilde{T}_{i}^{3}\rangle/\langle\tilde{T}_{i}^{2}\rangle^{3/2}$, which is shown in panel (f).
@@ 3.1 Construction of the Model Dynamical System
Such modes correspond to rigidly toroidally rotating fluctuations, with a toroidal angular phase velocity of $\Omega=\omega/n$. Hamiltonians consisting of rigidly toroidally rotating modes with a common angular phase velocity $\Omega$ will satisfy $\partial_{t}H+\Omega\partial_{\varphi}H=0$. Particles undergoing motion in such fields will have two invariants of motion, $(\mu,K)$, where $K=H-\Omega P_{\varphi}$ is the Hamiltonian in the rotating frame. This fact is often utilized in studies of energetic particle transport (Hsu and Sigmar, 1992; Todo, 2019). Note that a regime where most fluctuations share a nearly common angular phase velocity of $\Omega$ would correspond to weakly dispersive turbulence.
The key hypothesis we take is that in the presence of a zonal jet, the phenomenon of wave trapping will produce fluctuations which lead to this weakly dispersive regime. In Zhu et al. (2020a, b), it was demonstrated how wave trapping and anti-trapping play a key role determining the Dimits shift regime in fluid models of drift-wave turbulence. The key physics of wave trapping is that waves can constructively interfere. Physically, one can imagine that the $E\times B$ jet acts as a “cavity” which modifies the structure of the primary ITG instability. In the presence of shear of opposing signs, discrete quantum harmonic oscillator-like eigenmodes can appear which have radially-localized fluctuation envelopes. Intuitively, this radial localization increases the effective radial wavenumber, typically leading to a stabilizing effect on radial-gradient-driven instabilities. In such a regime turbulent mixing can become strongly spatially inhomogeneous; see for example Cao and Qi (2023, 2024) for detailed studies of the nonlinear dynamics of large-amplitude tertiary instabilities in a variant of the Hasegawa-Wakatani equations.
Figure 3: (a) Toroidal mode spectra and (b) angular phase velocities, both measured at one instant in time. $\langle\phi_{n}^{2}\rangle$ is the flux surface averaged squared electrostatic potential, and $\hat{\Omega}_{n}$ is the toroidally-directed angular phase velocity. The quantity $\Gamma_{0}k_{\theta}^{2}\langle\phi_{n}^{2}\rangle$ mimics the gyroaveraged $\langle\mathcal{J}[E_{\perp}]^{2}\rangle$ spectrum. Different toroidal mode numbers are shown with different colors and linestyles. $\Omega_{E}$ is shown on both plots as a solid purple line for reference, and the location of the shearless region in the zonal jet is demarcated in panel (a) by a dashed vertical gray line. A dashed horizontal black line is also plotted in in panel (b) over the angular phase velocities to show the rotation rate $\Omega$ used for the model fluctuations.
@@
Figure 4: Poloidal slices of (a) The electrostatic potential $\phi_{n}$ for the $n=39$ Fourier mode, and (b) the single-mode model electrostatic potential $\delta\hat{\phi}$ at a single instant in time. The fields are plotted against the radial coordinate $\psi_{n}$ and the perpendicular field line label $\alpha=\varphi-q\theta$. The shearless region $\psi=\psi^{*}$ is demarcated with a dashed gray line. (c) Space-time plot showing the value of the electrostatic potential fluctuations $\delta\phi$, evaluated at the shearless region $\psi=\psi^{*}$ plotted against $\alpha$. A line with $\varphi\propto\Omega t$ is shown demonstrating the fixed phase velocity of the fluctuations.